OpenStudy (anonymous):

cos theta(cos theta+3)=-2 solve for 180

OpenStudy (anonymous):

This is what I have: \[\cos \Theta(\cos \Theta + 3) = -2 \] Do you mean Solve when Theta is 180 Degrees?

OpenStudy (anonymous):

And is what I have written correct?

OpenStudy (anonymous):

yes that is correct

OpenStudy (raden):

i think the question asking find the theta in interval 0 <= theta <= 180

OpenStudy (anonymous):

Not to jump in on someone elses convo but i like really need help with this.. is anyone on this thing good at geometry ?

OpenStudy (anonymous):

yeah the question says "solve for \[\Theta if 0\le \Theta \le 180\]

OpenStudy (anonymous):

for \[\cos \Theta \left( \cos \theta +3 \right)=-2\]